A piece of copper is surrounded by air at a pressure of 1.01 x 105 Pa. The copper is placed into a chamber where the pressure is increased to 2.02 x 105 Pa. Determine the fractional change ??/K) in volume of the copper. The bulk modulus of copper is 1.3x 101 N/m2
Answer #1

The relation which relates pressure change with the fractional change in volume is

B =\frac{\Delta P}{\Delta V/V}

since bulk modules , B=1.3 *10^11 N/m^2

and P1=1.01 *10^5 pa, P2 =2.02*10^5 pa

So, \Delta P =P_2 -P_1 =2.02 \times 10^5 -1.01 \times 10^5 =1.01 \times 10^5 Pa

now we can write,

\Delta V/V =\frac{\Delta P}{B} =\frac{1.01 \times 10^5}{1.3 \times 10^{11}}=7.769 \times 10^{-7}

If you have any doubt in the solution then you can ask in the comment section.

good luck